the length of the side of a square is 3meters less than the length of the side of another square. The difference of their areas is 117 square meters. find the length of a side of each square.
Answer by scott8148(5891)

let s equal the smaller square ... (s+3)^2-s^2=117 ... s^2+6s+9-s^2=117 ... 6s=108 ... s=18 ... s+3=21